OpenStudy (anonymous):

I need to figure out how to solve for the compound inequality of -2(x+3)>9+3x

OpenStudy (texaschic101):

-2(x + 3) > 9 + 3x -- distribute through the parenthesis -2x - 6 > 9 + 3x -- subtract 3x from both sides -2x - 3x - 6 > 9 -- add 6 to both sides -2x - 3x > 9 + 6 -- combine like terms -6x > 15 -- divide by -6 and change the sign x < - 15/6

OpenStudy (texaschic101):

however, that is not a compound inequality

OpenStudy (anonymous):

Thats what I was thinking. I didnt think it was either. Its not for me actually its for my ex girlfriend. I was trying to tell her her teacher was wrong

OpenStudy (anonymous):

How did you get the 6 in the second part?

OpenStudy (texaschic101):

when you distribute through the parenthesis...that means you multiply the number outside of the parenthesis, by everything in the parenthesis... In other words..-2(x + 3) = -2 * x + (-2 * 3) = -2x - 6

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

Ohhh ok thank you so much

OpenStudy (texaschic101):

anytime :)